Do a two-sample test for equality of means assuming unequal variances. Calculate the p-value using Excel. (a-1) Comparison of GPA for randomly chosen college juniors and seniors: x⎯⎯1 = 4, s1 = .20, n1 = 15, x⎯⎯2 = 4.25, s2 = .30, n2 = 15, α = .025, left-tailed test. (Negative values should be indicated by a minus sign. Round down your d.f. answer to the nearest whole number and other answers to 4 decimal places. Do not use "quick" rules for degrees of freedom.) d.f. t-calculated p-value .0088 t-critical -2.145 (a-2) Based on the above data choose the correct decision. Reject the null hypothesis Do not reject the null hypothesis (b-1) Comparison of average commute miles for randomly chosen students at two community colleges: x⎯⎯1 = 17, s1 = 5, n1 = 22, x⎯⎯2 = 21, s2 = 7, n2 = 19, α = .05, two-tailed test. (Negative values should be indicated by a minus sign. Round down your d.f. answer to the nearest whole number and other answers to 4 decimal places. Do not use "quick" rules for degrees of freedom.) d.f. t-calculated p-value t-critical +/- (b-2) Based on the above data choose the correct decision. Do not reject the null hypothesis Reject the null hypothesis (c-1) Comparison of credits at time of graduation for randomly chosen accounting and economics students: x⎯⎯1 = 141, s1 = 2.8, n1 = 12, x⎯⎯2 = 138, s2 = 2.7, n2 = 17, α = .05, right-tailed test. (Negative values should be indicated by a minus sign. Round down your d.f. answer to the nearest whole number and other answers to 4 decimal places. Do not use "quick" rules for degrees of freedom.) d.f. t-calculated p-value t-critical (c-2) Based on the above data choose the correct decision. Reject the null hypothesis Do not reject the null hypothesis
Answer #1

Sol:

Part a)

Test Statistic :-
(X-X)/(S/n1)+(S3/n2)
t = (4.05 – 4.35)/V(0.22/15) + (0.32/15)
t = -3.2225


Test Criteria :-
Reject null hypothesis if t < -t(α, DF)
DF = ((Si/nl + S3/n2))/((Si/nl)?/n1 – 1) + (S/n2)²/n2 – 1))
((0.2/15 0.3/15))/((0.2/15/15 1)(0.32/15)2/15 1)) DF
DF = 24


Critical value t(α, DF) = t( 0.025 , 24 ) = 2.064
t < -t(α, DF) = -3.2225 < -2.064
Result :- Reject Null Hypothesis

Decision based on P value

P - value = P ( t > 3.2225 ) = 0.0018
Reject null hypothesis if P value < α level of significance
P - value = 0.0018 < 0.025 ,hence we reject null hypothesis
Conclusion :- Reject null hypothesis

Part b)

Test Statistic :-
(X-X)/(S/n1)+(S3/n2)
t = (19 – 25)/(52/22) + (72/19)
t = -3.1128


Test Criteria :-
Reject null hypothesis if | t | > t(α/2, DF)
DF = ((Si/nl + S3/n2))/((Si/nl)?/n1 – 1) + (S/n2)²/n2 – 1))
DF = ((52/22 + 72/19))/((52/22)/22 - 1) + (72/19)2/19 – 1))
DF = 32

Critical value   t(α/2, DF) = t(0.05 /2, 32 ) = 2.037

| t | > t(α/2, DF) = 3.1128 > 2.037
Result :- Reject Null Hypothesis


Decision based on P value
P - value = P ( t > 3.1128 ) = 0.0039
Reject null hypothesis if P value < α = 0.05 level of significance
P - value = 0.0039 < 0.05 ,hence we reject null hypothesis
Conclusion :- Reject null hypothesis

Part c)

Test Statistic :-
(X-X)/(S/n1)+(S3/n2)
t = (144 – 143)/V (2.82/12) + (2.72/17)
t = 0.9613

Test Criteria :-
Reject null hypothesis if t > t(α, DF)
DF = ((Si/nl + S3/n2))/((Si/nl)?/n1 – 1) + (S/n2)²/n2 – 1))
DF = ((2.82/12 +2.72/17))?/((2.84/12²/12 – 1) + (2.72/17)?/17 – 1))
DF = 23


Critical value   t(α, DF) = t( 0.05 , 23 ) = 1.714

t > t(α, DF) = 0.9613 < 1.714
Result :- Fail to Reject Null Hypothesis


Decision based on P value
P - value = P ( t > 0.9613 ) = 0.1732
Reject null hypothesis if P value < α level of significance
P - value = 0.1732 > 0.05 ,hence we fail to reject null hypothesis
Conclusion :- We Accept H0